OpenStudy (anonymous):

2x+1)/(4)+(4x+1)/(6)=(1)

OpenStudy (anonymous):

ah solve for x make your life easy and multiply both sides by 12 to clear the fractions

OpenStudy (anonymous):

\[3(2x+1)+2(4x+1)=12\] and then it will be much easier

OpenStudy (anonymous):

multiply out, combine like terms etc. you good from there?

OpenStudy (anonymous):

urm no tottally sure but this is wat I got wen i tried it..

OpenStudy (anonymous):

did you get \[x=\frac{1}{2}\]?

OpenStudy (anonymous):

no i got 6x/6 =22/6

OpenStudy (anonymous):

which gave me x=3.666667

OpenStudy (anonymous):

=/ dnt think its right is it

OpenStudy (anonymous):

start here \[3(2x+1)+2(4x+1)=12\]

OpenStudy (anonymous):

how did you get 3(2x+1)+2(4x+1)=12

OpenStudy (anonymous):

then multiply out using the distributive property \[6x+3+8x+2=12\] combine like terms \[14x+5=12\] subtract 5 \[14x=7\] divide by 14 \[x=\frac{1}{2}\]

OpenStudy (anonymous):

im confused =/ sorry...

OpenStudy (anonymous):

multiply both sides by the least common multiply of 4 and 6, which is 12 i will write the steps

sam (.sam.):

(2x+1)/(4)+(4x+1)/(6)=(1) (2x+1)/(4) * (3)/(3) + (4x+1)/(6) * (2)/(2)=(1) (2x+1)(3)/(12)+((4x+1)(2))/(12)=(1) ((2x+1)(3)+((4x+1)(2)))/(12)=(1) (14x+5)/(12)=(1) 14x+5=12 14x=7 x=1/2

OpenStudy (anonymous):

\[\frac{2x+1}{4}+\frac{4x+1}{6}=1\] \[12\left(\frac{2x+1}{4}+\frac{4x+1}{6}\right)=1\times 12\] \[3(2x+1)+2(4x+1)=12\] the last line because 4 goes in to 12 3 times and 6 goes in to 12 2 times

OpenStudy (anonymous):

don't forget the parentheses because you have to mulitply out using the distributive law

OpenStudy (anonymous):

okay thanks.